# Integrating Personality into Digital Humans: A Review of LLM-Driven Approaches for Virtual Reality

### v1

@@ 3.1 Qualitative Evaluation
As previously discussed, evaluating the personality traits of digital humans involves complex, nuanced, and non-standardized methods. Among these, qualitative evaluation methods are widely used, particularly in LLM-based conversational models [45]. These methods are primarily subjective, relying on human (or, in some cases, LLM) interpretation to assess an agent’s performance, coherence, and adherence to specific personality traits. While qualitative evaluations are valuable for understanding how users perceive an agent’s behavior and personality—critical for improving user-agent relationships—they lack standardization [9]. To address the limitations of purely human or purely LLM-based evaluations, many studies adopt a hybrid approach, combining the strengths of both methods [35]. However, the inherent subjectivity of these methods means that experiences and interpretations can vary significantly from person to person (or model to model, in the case of LLM-as-judge evaluations).
@@
Human evaluation has historically been one of the most widely used qualitative methods in research, particularly for assessing alignment in LLM-based models. This approach evaluates whether an LLM’s behavior aligns with human intentions, ethical values, and predefined goals [1]. Beyond alignment, it serves as a critical tool for evaluating the personality traits and behavioral consistency of LLMs [82]. This method typically involves interactive tests, where evaluators engage directly with the agent to determine whether it effectively exhibits the desired personality traits and meets user expectations [35]. Human evaluation can be conducted from multiple perspectives, enabling a comprehensive analysis of various dimensions of the agent’s behavior. These distinct angles of evaluation provide deeper insights into the agent’s overall performance and its adherence to intended personality traits and behavioral standards.
@@
Human evaluation has been widely used to assess how personality traits influence user perceptions of virtual agents during conversational tasks, offering valuable insights into social interactions. Research in this area has explored virtual agents designed to exhibit extroverted or introverted characteristics, expressed through variations in tone, communicative style, and level of engagement [40]. User assessments based on factors such as warmth, empathy, and realism have demonstrated the significant role of personality traits in shaping perceptions and interactions in virtual environments.
@@
However, human evaluation as a methodology presents notable limitations. Its subjective nature introduces variability, as individual evaluators may interpret and score interactions differently, leading to inconsistencies, particularly in studies involving multiple evaluators with diverse perspectives. This variability can undermine the reliability of results and complicate cross-study comparisons. Furthermore, the effectiveness of human evaluation as a primary approach to assess the personality traits of digital humans remains a topic of ongoing debate within the research community [13]. These challenges highlight the need for standardized evaluation protocols and objective metrics to improve consistency and comparability between studies.
@@
Another qualitative evaluation method involves the use of the LLM-as-judge approach. This emerging framework employs LLMs to evaluate the personality and behavior of other LLMs and conversational agents by simulating human judgment [45]. The LLM-as-judge methodology has gained traction due to its scalability and ability to automate the evaluation process [17, 97], reducing the dependence on human evaluators, which can be time-consuming and costly.
@@
In this approach, an LLM is provided with a prompt containing the text to be assessed along with clear evaluation criteria. For example, the criteria might involve a question designed to assess specific personality traits of the tested LLM. Based on the prompt, the evaluating LLM assigns scores or labels to the text according to the predefined criteria. This method has been widely employed to evaluate conversational agents, models, and chatbots across various domains [84], including the context of this research [80].
@@
Although the LLM-as-judge approach offers notable benefits, such as increased efficiency and reduced costs for evaluating LLM personality traits, it presents several challenges, including inherent biases in the models, sensitivity to different prompts, limitations specific to individual models, and concerns regarding the consistency and reliability of the evaluations [10, 95]. Addressing these issues is essential to enhance the accuracy and fairness of this methodology, highlighting the need for further refinement and the establishment of standardized evaluation frameworks.
@@ 3.2 Quantitative Evaluation
Common approaches, such as personality assesments [86] and word frequency analysis [74], aim to provide standardization and reproducibility, addressing the variability and subjectivity often associated with qualitative assessments. However, they still face challenges, including ongoing debates about their reliability in accurately evaluating LLM personality traits [24]. Despite these limitations, quantitative methods provide a valuable foundation for comparing models and identifying consistent patterns in personality expression, making them indispensable for evaluating digital agents in a systematic manner.
@@
Among quantitative evaluation methods, personality assessments originally developed for the psychological evaluation of humans [62, 86] have been extensively employed to analyze a wide range of personality characteristics in LLM-driven agents. These assessments serve as systematic tools designed to measure diverse dimensions of personality, encompassing traits, behaviors, motivations, and interpersonal dynamics. Instruments such as the Big Five Inventory (BFI) [36], the International Personality Item Pool (IPIP) [22], and the NEO Personality Inventory (NEO-PI-3) [51] are among the most frequently utilized for evaluating both human subjects and virtual agents.
@@
In these structured assessments, individuals or LLMs provide evaluations of behaviors, thoughts, and feelings by responding to a series of predetermined statements or questions. These responses, collected through formats such as Likert scales (e.g., ratings ranging from ”strongly disagree” to ”strongly agree”) [48], true-false options, or forced-choice formats, are scored and analyzed to produce detailed profiles of personality traits. When applied to LLMs, specific prompt instructions are often used to simulate self-reported responses, enabling the modeling of human-like personality traits [34, 67, 64].
@@
Although widely used, these assessments present significant challenges when applied to non-human entities such as LLMs. One notable issue is the sensitivity of LLMs to multiple-choice question (MCQ) formats, as responses are heavily influenced by the order of options presented [93]. Since self-reported questionnaires frequently rely on MCQs, this introduces bias into the responses generated by digital agents, compromising the reliability of the evaluations. Consequently, the administration of self-reported questionnaires to virtual agents often yields unreliable results [24].
@@
To address these challenges, efforts have been made to adapt self-report questionnaires originally designed for humans into formats suitable for assessing personality traits in digital agents [42]. One significant contribution is the Language Model Linguistic Personality Assessment (LMLPA) [94]. This system adapts the Big Five Inventory — a cornerstone in psychometric research [36] — by aligning its format with the operational capabilities of LLMs. Incorporating AI raters and linguistic prompts, the LMLPA quantitatively evaluates personality traits through the linguistic outputs of LLMs, offering a novel, domain-specific approach. Although promising, the methodology requires further empirical validation to ensure its robustness and applicability across diverse scenarios [94].
@@
Another widely employed quantitative evaluation method for assessing personality traits in LLM-based digital agents is the Linguistic Inquiry and Word Count (LIWC) software [59]. LIWC analyzes text by mapping words and phrases to a curated dictionary, categorizing them into dimensions reflecting psychological, emotional, and social constructs [74]. Its latest version, LIWC-22, includes over 12,000 words and expressions across 117 categories, such as personal pronouns, emotion-related vocabulary, and cognitive indicators [3].
@@
Widely applied in psychology [78], and communication research [26], LIWC is a valuable tool for studying personality traits in digital agents [74]. In LLM-based digital agents, LIWC has been applied to assess alignment [21], personality retention, and emotional embedding [79]. These applications highlight LIWC’s ability to reveal subtle patterns and provide meaningful insights into how LLMs convey and replicate personality traits through language.
@@
While qualitative methods are effective for analyzing personality traits in text-based agents, they are less suited for evaluating digital humans in VR environments. These agents express personality through a combination of verbal outputs and non-verbal behaviors, such as gestures, facial expressions, and vocal tone. This multimodal nature introduces complexities that demand updated evaluation methods to fully capture their dynamics. Embodied conversational agents, for example, offer a heightened sense of social presence and interaction, resembling face-to-face communication [70]. Solely relying on textual analysis neglects critical non-verbal cues, such as gestures and intonation, which can either reinforce or contradict verbal outputs.
@@ 4 Future
Despite rapid advancements in the field of digital humans, the use of cutting-edge AI technologies, particularly LLMs, to simulate human identity and personality in realistic, non-chat-based scenarios remains underexplored. The capabilities of generative models to mimic human behavior present a promising avenue for applications across diverse domains. For instance, LLMs can enable adaptive personalities in NPC characters, allowing non-scripted, context-driven conversations. These NPCs could maintain both short-term and long-term memory during interactions, enriching gaming experiences with more dynamic and immersive storytelling [96]. Beyond entertainment, educational applications of such agents have shown potential in improving learning outcomes, such as enhancing user satisfaction and reducing the time required to master new concepts [23, 75].
An emerging research area involves leveraging the generative capabilities of LLMs to produce text that not only conveys conversational content but also encodes directives for corresponding visual actions. By integrating emotion recognition and generation, these systems could embed instructions for digital humans to perform specific facial expressions or body movements aligned with the intended emotional tone of the interaction. This approach enables the generated text to act as both a communication medium and a control mechanism for non-verbal expressions, enhancing the realism and interactivity of virtual environments [28]. Such multimodal integration bridges the gap between verbal and non-verbal communication, creating a more seamless and human-like interaction experience.
However, while LLMs exhibit remarkable capacities for simulating and adapting human personalities, they are hindered by significant computational demands. The high inference time associated with auto-regressive generation methods translates into latency challenges, requiring substantial GPU infrastructure to ensure real-time responses [90, 43]. On the other hand, recent research on small LLMs (SLLMs) suggests that these models can perform tasks previously reserved for larger models [11]. Although still in its initial research, this field holds potential for addressing the computational challenges of LLMs, enabling the development of lightweight systems capable of delivering personality simulation and emotion-driven expressions in real time. Further research is needed to evaluate the effectiveness of SLLMs in these complex tasks, particularly in scenarios requiring both verbal and non-verbal adaptability.

### v2

@@ 3.1 Qualitative Evaluation
As previously discussed, evaluating the personality traits of digital humans involves complex, nuanced, and non-standardized methods. Among these, qualitative evaluation methods are widely used, particularly in LLM-based conversational models [li2024llmsasjudgescomprehensivesurveyllmbased]. These methods are primarily subjective, relying on human (or, in some cases, LLM) interpretation to assess an agent’s performance, coherence, and adherence to specific personality traits. While qualitative evaluations are valuable for understanding how users perceive an agent’s behavior and personality—critical for improving user-agent relationships—they lack standardization [chen-etal-2024-humans]. To address the limitations of purely human or purely LLM-based evaluations, many studies adopt a hybrid approach, combining the strengths of both methods [jiang-etal-2024-personallm]. However, the inherent subjectivity of these methods means that experiences and interpretations can vary significantly from person to person (or model to model, in the case of LLM-as-judge evaluations).
@@
Human evaluation has historically been one of the most widely used qualitative methods in research, particularly for assessing alignment in LLM-based models. This approach evaluates whether an LLM’s behavior aligns with human intentions, ethical values, and predefined goals [abeysinghe2024challengesevaluatingllmapplications]. Beyond alignment, it serves as a critical tool for evaluating the personality traits and behavioral consistency of LLMs [vu2025psychadapteradaptingllmtransformers]. This method typically involves interactive tests, where evaluators engage directly with the agent to determine whether it effectively exhibits the desired personality traits and meets user expectations [jiang-etal-2024-personallm]. Human evaluation can be conducted from multiple perspectives, enabling a comprehensive analysis of various dimensions of the agent’s behavior. These distinct angles of evaluation provide deeper insights into the agent’s overall performance and its adherence to intended personality traits and behavioral standards.
@@
Human evaluation has been widely used to assess how personality traits influence user perceptions of virtual agents during conversational tasks, offering valuable insights into social interactions. Research in this area has explored virtual agents designed to exhibit extroverted or introverted characteristics, expressed through variations in tone, communicative style, and level of engagement [kroczek2024influencepersonaconversationaltask]. User assessments based on factors such as warmth, empathy, and realism have demonstrated the significant role of personality traits in shaping perceptions and interactions in virtual environments.
@@
However, human evaluation as a methodology presents notable limitations. Its subjective nature introduces variability, as individual evaluators may interpret and score interactions differently, leading to inconsistencies, particularly in studies involving multiple evaluators with diverse perspectives. This variability can undermine the reliability of results and complicate cross-study comparisons. Furthermore, the effectiveness of human evaluation as a primary approach to assess the personality traits of digital humans remains a topic of ongoing debate within the research community [clark2021thatshumangoldevaluating]. These challenges highlight the need for standardized evaluation protocols and objective metrics to improve consistency and comparability between studies.
@@
Another qualitative evaluation method involves the use of the LLM-as-judge approach. This emerging framework employs LLMs to evaluate the personality and behavior of other LLMs and conversational agents by simulating human judgment [li2024llmsasjudgescomprehensivesurveyllmbased]. The LLM-as-judge methodology has gained traction due to its scalability and ability to automate the evaluation process [dong2024llmpersonalizedjudge, zhu2023judgelmfinetunedlargelanguage], reducing the dependence on human evaluators, which can be time-consuming and costly.
@@
In this approach, an LLM is provided with a prompt containing the text to be assessed along with clear evaluation criteria. For example, the criteria might involve a question designed to assess specific personality traits of the tested LLM. Based on the prompt, the evaluating LLM assigns scores or labels to the text according to the predefined criteria. This method has been widely employed to evaluate conversational agents, models, and chatbots across various domains [wei2024systematicevaluationllmasajudgellm], including the context of this research [tseng-etal-2024-two].
@@
Although the LLM-as-judge approach offers notable benefits, such as increased efficiency and reduced costs for evaluating LLM personality traits, it presents several challenges, including inherent biases in the models, sensitivity to different prompts, limitations specific to individual models, and concerns regarding the consistency and reliability of the evaluations [chen2024humansllmsjudgestudy, zheng2023judgingllmasajudgemtbenchchatbot]. Addressing these issues is essential to enhance the accuracy and fairness of this methodology, highlighting the need for further refinement and the establishment of standardized evaluation frameworks.
@@ 3.2 Quantitative Evaluation
Common approaches, such as personality assesments [weiner2017handbook] and word frequency analysis [tausczik2010psychological], aim to provide standardization and reproducibility, addressing the variability and subjectivity often associated with qualitative assessments. However, they still face challenges, including ongoing debates about their reliability in accurately evaluating LLM personality traits [gupta-etal-2024-self]. Despite these limitations, quantitative methods provide a valuable foundation for comparing models and identifying consistent patterns in personality expression, making them indispensable for evaluating digital agents in a systematic manner.
@@
Among quantitative evaluation methods, personality assessments originally developed for the psychological evaluation of humans [pilbeam2012psychometric, weiner2017handbook] have been extensively employed to analyze a wide range of personality characteristics in LLM-driven agents. These assessments serve as systematic tools designed to measure diverse dimensions of personality, encompassing traits, behaviors, motivations, and interpersonal dynamics. Instruments such as the Big Five Inventory (BFI) [john1991big], the International Personality Item Pool (IPIP) [goldberg2006international], and the NEO Personality Inventory (NEO-PI-3) [mccrae2005neo] are among the most frequently utilized for evaluating both human subjects and virtual agents.
@@
In these structured assessments, individuals or LLMs provide evaluations of behaviors, thoughts, and feelings by responding to a series of predetermined statements or questions. These responses, collected through formats such as Likert scales (e.g., ratings ranging from ”strongly disagree” to ”strongly agree”) [likert1932technique], true-false options, or forced-choice formats, are scored and analyzed to produce detailed profiles of personality traits. When applied to LLMs, specific prompt instructions are often used to simulate self-reported responses, enabling the modeling of human-like personality traits [jiang_2024, serapiogarcía2023personalitytraitslargelanguage, rao-etal-2023-chatgpt].
@@
Although widely used, these assessments present significant challenges when applied to non-human entities such as LLMs. One notable issue is the sensitivity of LLMs to multiple-choice question (MCQ) formats, as responses are heavily influenced by the order of options presented [zheng2024largelanguagemodelsrobust]. Since self-reported questionnaires frequently rely on MCQs, this introduces bias into the responses generated by digital agents, compromising the reliability of the evaluations. Consequently, the administration of self-reported questionnaires to virtual agents often yields unreliable results [gupta-etal-2024-self].
@@
To address these challenges, efforts have been made to adapt self-report questionnaires originally designed for humans into formats suitable for assessing personality traits in digital agents [lee2024llmsdistinctconsistentpersonality]. One significant contribution is the Language Model Linguistic Personality Assessment (LMLPA) [zheng2024lmlpalanguagemodellinguistic]. This system adapts the Big Five Inventory — a cornerstone in psychometric research [john1991big] — by aligning its format with the operational capabilities of LLMs. Incorporating AI raters and linguistic prompts, the LMLPA quantitatively evaluates personality traits through the linguistic outputs of LLMs, offering a novel, domain-specific approach. Although promising, the methodology requires further empirical validation to ensure its robustness and applicability across diverse scenarios [zheng2024lmlpalanguagemodellinguistic].
@@
Another widely employed quantitative evaluation method for assessing personality traits in LLM-based digital agents is the Linguistic Inquiry and Word Count (LIWC) software [pennebaker2001linguistic]. LIWC analyzes text by mapping words and phrases to a curated dictionary, categorizing them into dimensions reflecting psychological, emotional, and social constructs [tausczik2010psychological]. Its latest version, LIWC-22, includes over 12,000 words and expressions across 117 categories, such as personal pronouns, emotion-related vocabulary, and cognitive indicators [boyd2022development].
@@
Widely applied in psychology [tov2013detecting], and communication research [haber2015stabilityonlinelanguagefeatures], LIWC is a valuable tool for studying personality traits in digital agents [tausczik2010psychological]. In LLM-based digital agents, LIWC has been applied to assess alignment [frisch2024llmagentsinteractionmeasuring], personality retention, and emotional embedding [tripto2024shiptheseuscuriouscases]. These applications highlight LIWC’s ability to reveal subtle patterns and provide meaningful insights into how LLMs convey and replicate personality traits through language.
@@
While qualitative methods are effective for analyzing personality traits in text-based agents, they are less suited for evaluating digital humans in VR environments. These agents express personality through a combination of verbal outputs and non-verbal behaviors, such as gestures, facial expressions, and vocal tone. This multimodal nature introduces complexities that demand updated evaluation methods to fully capture their dynamics. Embodied conversational agents, for example, offer a heightened sense of social presence and interaction, resembling face-to-face communication [10.1145/3173574.3173863]. Solely relying on textual analysis neglects critical non-verbal cues, such as gestures and intonation, which can either reinforce or contradict verbal outputs.
@@ 4 Future
Despite rapid advancements in the field of digital humans, the use of cutting-edge AI technologies, particularly LLMs, to simulate human identity and personality in realistic, non-chat-based scenarios remains underexplored. The capabilities of generative models to mimic human behavior present a promising avenue for applications across diverse domains. For instance, LLMs can enable adaptive personalities in NPC characters, allowing non-scripted, context-driven conversations. These NPCs could maintain both short-term and long-term memory during interactions, enriching gaming experiences with more dynamic and immersive storytelling [zheng2024memorynpc]. Beyond entertainment, educational applications of such agents have shown potential in improving learning outcomes, such as enhancing user satisfaction and reducing the time required to master new concepts [pedagogic_agent, taylor2023rise].
An emerging research area involves leveraging the generative capabilities of LLMs to produce text that not only conveys conversational content but also encodes directives for corresponding visual actions. By integrating emotion recognition and generation, these systems could embed instructions for digital humans to perform specific facial expressions or body movements aligned with the intended emotional tone of the interaction. This approach enables the generated text to act as both a communication medium and a control mechanism for non-verbal expressions, enhancing the realism and interactivity of virtual environments [hasan_sapienaffective]. Such multimodal integration bridges the gap between verbal and non-verbal communication, creating a more seamless and human-like interaction experience.
However, while LLMs exhibit remarkable capacities for simulating and adapting human personalities, they are hindered by significant computational demands. The high inference time associated with auto-regressive generation methods translates into latency challenges, requiring substantial GPU infrastructure to ensure real-time responses [xia2023flashLLM, li2024llminferenceservingsurvey]. On the other hand, recent research on small LLMs (SLLMs) suggests that these models can perform tasks previously reserved for larger models [chen2024rolesmallmodelsllm]. Although still in its initial research, this field holds potential for addressing the computational challenges of LLMs, enabling the development of lightweight systems capable of delivering personality simulation and emotion-driven expressions in real time. Further research is needed to evaluate the effectiveness of SLLMs in these complex tasks, particularly in scenarios requiring both verbal and non-verbal adaptability.